Exp. 9 Evaluating Antacids A roll of 12 brand A tablets costs so.59. How many moles of hydrochloric acid will a dollars worth of brand A tablets neutralize? Just as in the experiment you will do, the second half of the tablet was evaluated using the same process. However, in breaking the tablet in two, some crumbs of the tablet were left on the bench top, and not included in either titration flask. Will this make any difference in the final results. Explain. 2. Brand Bantacid sells 100 tablets in a bottle for $3.99. A student found that a brand B tablet would neutralize 0.011 moles of HCI per tablet. Which is the better buy brand A or brand B? Explain. 3. In brand c, each tablet contains 550mg of calcium carbonate and 110mg of magnesium hydroxide.
media%2F54e%2F54ec3eba-079e-41d4-814f-d9
Answer #1

1. first determine the moles of the HCL was added.

moles = molarity * liters

we consider here molarity is = 1.00 & liters is also=1.0 lit

so, moles = 1 mole

so the moles of NaOH required to neutralize is = 1.0

HCL + NaOH ---> NaCl + H2O

as per above equation 1 mole of HCl reacts with 1 mole of NaOH to produce 1 mole of NaCl & 1 mole of H2O.

so, in brand B 0.011 mole of Hcl converted to 0.011 mole of NaCl anf H2O.

so, Total number of moles of HCl required to neutralize the antacid tablet = 1.0-0.011 = 0.989 moles

so, brand B is better to buy.

2. now in brand C 660 mg = 0.66 gram of tablet contains 0.989 moles of antacids.

according to B is in 100 tablets so, cost per tablet is = 1/100 = $0.0100

   cost per mole of antacid = $0.01 / 0.989 = $0.0101